Two reactions and their equilibrium constants are given.

A+2B−⇀2C K1=2.93

2C−⇀D K2=0.258

Calculate the value of the equilibrium constant for the reaction

D−>A+2B
.
1
Expert's answer
2021-04-12T03:46:27-0400

Adding the two equations gives:

A + 2B = 2C K1=2.93

+

2C = D K2=0.258

Result:

A + 2B + 2C 2C + D

K1 + K2

Remove 2С:

A + 2B = D

K1 + K2 = 2.93 + 0.258 = 3.188

Reversed reaction:

D = A + 2B

K3 = 1 / 3.188 = 0.314